# Transfer-based Adversarial Poisoning Attacks for Online (MIMO-)Deep Receviers

### v2

@@ V-C Online Training Methods
The objective of this paper is to present an attack strategy, which attacks online adaptation of the deep receivers. Consequently, the focus is on the deep receiver’s online training, with the parameter configurations illustrated in Table I. The parameter settings for online training come from [15, 17]. Upon receipt of a data block, the deep receiver is only able to utilize a subset of the data block for training, like pilot symbols, specifically $L_{\text{pilot}}=200$. Then the deep reciver predicts the subsequent $L_{\text{info}}=50000$ symbols. In the experiment, a total of 100 data blocks were transmitted, with the transmitted data being QPSK modulated, i.e., the user-transmitted symbols $\mathbf{s}$ were mapped to the set $C=\left\{\left(\pm\frac{1}{\sqrt{2}},\pm\frac{1}{\sqrt{2}}\right)\right\}^{4}$. Furthermore, the deep receivers are trained using the Adam optimizer[51]. The training epochs are 300. The initial learning rate $\eta$ is set to 5 ×$\times 10^{-3}$ for ResNet detector, black-box DNN detector and DeepSIC. Additionally, online training is performed for different architectural receivers, which were implemented in the following two cases:
@@
    Online learning: Based on the adaptation to the pilot data from the previous data block, the deep receiver trains and updates the current model using the limited pilot symbols received in the current data block.
@@
    Online meta-learning: According to [20], [52], the meta-learning framework is employed to facilitate adaptation to the channel. In particular, the pilot data from 5 data blocks is accumulated each time, after which meta-learning is performed on the accumulated data in order to obtain the meta-learning weights of the deep receiver. Subsequently, the aforementioned meta-learning weights are employed in an online learning process involving the pilot data of the current block.
Unless otherwise stated, the deep receivers are trained using online learning, including black-box DNN detector, ResNet detector, and DeepSIC. Only online meta-learning DeepSIC receiver, i.e., Meta-DeepSIC, is trained using online meta-learning methods.
@@ V-D Attacker’s Configuration
The attack samples are designed based on the gradient of the surrogate model, and joint learning is performed on the surrogate model based on the collected pilot dataset. In the joint learning configuration, the black-box DNN detector, presented in Section V-B, is employed as the surrogate model. This model has three times the number of parameters compared to a single sub-network of DeepSIC. The parameter configurations for joint learning are referenced in [15]. In particular, a linear time-varying synthetic channel model is employed to generate the channel data for joint learning. The training data is generated under the condition of an SNR of 2 dB to 16 dB with an interval of 2 dB, and the SNR used for generation is represented as $SNR_{\text{sur}}$. In this manner, $L_{\text{sur}}=5000$ training pilot symbols are generated at each SNR value, and the surrogate model is trained in accordance with this procedure.
Moreover, the poisoning samples are optimized iteratively using the PGD algorithm. In this context, the adversarial poisoning attack samples are created with reference to the settings in [44], and the adversarial poisoning samples at this specific parameter configuration are the most toxic. The iteration step size is set to $\gamma=0.01$, the iterations of PGD is $Q=250$, and the upper bound $\epsilon$ of the perturbation under the constrained norm $p=2$ is set to 0.3. The maximum and minimum values of the received symbol magnitude are denoted as $I_{max}$ and $I_{min}$, respectively, and $I_{max}=max\{\|\mathbf{y}\|\}=-I_{min}$. The attack samples are presented in Fig. 8 and Fig. 9, respectively. The two subfigures in Fig. 8 represent the real and imaginary parts of the original and poisoned received symbols, respectively. Fig. 9 presents the original and poisoned received symbols in a constellation diagram.
@@ V-E Numerical Results under Four Channel Models
This section presents the results of an experimental investigation into the efficacy of the proposed poisoning attack method in the context of various channel models. Specifically, the effectiveness of the method is evaluated under the following scenarios: Linear time-varying synthetic channels, nonlinear time-varying synthetic channels, linear static synthetic channels, and time-varying COST 2100 channels. For black-box DNN detector, a white-box poisoning attack is launched against this architecture receiver since its architecture is the same as the surrogate model. For the ResNet detector, we implement transfer poisoning attacks in black box scenarios. As the DeepSIC comprises multiple sub-networks, it is not feasible to utilize its gradient information directly in the design of poisoning attack samples. It is therefore anticipated that the attack perturbations designed on the surrogate model will transfer the poisoning effect to DeepSIC. Consequently, a transfer-based poisoning attack is executed on DeepSIC.
@@ V-E1 Linear Time-varying Synthesis Channel Results
Firstly, Fig. 10 evaluates the effectiveness of the proposed poisoning attack method under a linear time-varying synthetic channel.
Fig. 10(a) illustrates the results of 100 data transmission blocks, with the cumulative SER calculated on a block-wise basis over five repetitions at an SNR of 14 dB. It can be observed that the Meta-DeepSIC, which combines the model-based method and the meta-learning, is more effective in capturing the time-varying characteristics of the wireless channel in the absence of a poisoning attack. Consequently, it achieves a superior SER performance in comparison to the black-box DNN detector, ResNet detector and the DeepSIC. However, the Meta-DeepSIC is also more susceptible to poisoning attacks than the DeepSIC. This finding aligns with the conclusions presented in [46]. Furthermore, the proposed black box poisoning attack significantly impacts the ResNet detector. Since this method targets the black-box DNN detector through a white-box attack approach, the black-box DNN detector performs the worst after poisoning.
Fig. 10(b) illustrates the results of the 100-block data transmission experiment, with the SER averaged across the same five repetitions under varying SNR conditions. It can be observed that the proposed poisoning attack is effective for ResNet detector, black-box DNN detector and DeepSIC at different SNRs. In particular, the SER performance of the poisoned DeepSIC exhibits a degradation of approximately 0.67 dB compared to that of the normal DeepSIC at an SNR of 10 dB. Similarly, the SER performance of the poisoned Meta-DeepSIC is seen to deteriorate by approximately 0.91 dB in comparison to that of the normal Meta-DeepSIC. Additionally, it is observed that the SER of the black-box DNN detector exhibits an increase of approximately 0.5 dB, and the SER of the ResNet detector exhibits an increase of approximately 0.68 dB. As the SNR increases, the impact of the attack becomes increasingly pronounced. The deterioration in the SER performance of the poisoned DeepSIC and Meta-DeepSIC reaches 1.37 dB and 2.0 dB, respectively, at SNR = 15 dB.
@@ V-E2 Non-linear Time-varying Synthetic Channels Results
Secondly, Fig. 11 evaluates the effectiveness of the proposed poisoning attack methodology under a nonlinear time-varying synthetic channel. The experimental configuration are consistent with those depicted in Fig. 10. It can be observed that the proposed poisoning attack method results in a more pronounced poisoning effect in the nonlinear time-varying synthetic channel environment than in the linear time-varying synthetic channel. In particular, the proposed poisoning attack method has a detrimental impact on the SER performance of the ResNet detector, the black-box DNN detector, DeepSIC, and Meta-DeepSIC, with an SER performance deterioration of 0.93 dB, 0.53 dB, 1.35 dB, and 1.42 dB, respectively, at SNR = 10 dB. Similarly, as the SNR increases, the attack effect becomes more pronounced, with the SER performance deterioration of the four receiver architectures reaching 1.13 dB, 0.71 dB, 2.74 dB, and 3.09 dB, respectively, at SNR = 15 dB.
@@ V-E3 Linear Static Synthetic Channels Results
Fig. 12 presents an evaluation of the efficacy of the proposed poisoning attack method in the context of a linear static synthetic channel. Similarly, the experimental scenario and parameter configuration are consistent with those depicted in Fig. 10. It can be observed that the performance of DeepSIC and Meta-DeepSIC is essentially indistinguishable in this case. Furthermore, the lack of diverse data renders the black-box DNN detector unsuitable for adapting to the channel environment. Meanwhile, compared to the black box DNN detector, the ResNet detector that performed better in the first two channel environments performed the worst in the static channel model. The poisoning effect of the ResNet detector has lost its reference significance. In particular, in this scenario with an SNR of 10 dB, the proposed poisoning attack method results in an SER performance degradation of 0.08 dB, 0.71 dB, and 0.72 dB for the black-box architecture receiver, the DeepSIC receiver, and the Meta-DeepSIC receiver, respectively. As with the preceding scenario, the impact of the attack is more pronounced as the SNR increases. At an SNR of 15 dB, the SER performance degradation for the three receiver architectures is 0.06 dB, 1.05 dB, and 1.07 dB, respectively.
@@ V-E4 Time-varying COST 2100 Channels Results
The efficacy of the proposed poisoning attack method is subsequently assessed under the time-varying COST 2100 channel, with the experimental results illustrated in Fig. 13. The experimental configuration is identical to that of the linear time-varying synthetic channel. Concurrently, the surrogate model is trained on the joint channel data based on the time-varying linear channel model. As illustrated in the Fig. 13, the proposed poisoning attack method is not only effective for receivers with different architectural depths and different online training scenario settings, but also adaptable to different channel environments. In this channel environment and with an SNR of 10 dB, the poisoning attack results in a deterioration of SER performance for the ResNet detector, the black-box DNN detector, DeepSIC, and Meta-DeepSIC of up to 0.12 dB, 0.25 dB, 0.61 dB, and 0.78 dB, respectively. At an SNR of 15 dB, the deterioration of SER performance for the four receiver architectures is up to 0.04 dB, 0.23 dB, 2.11 dB, and 1.36 dB, respectively.
@@ V-E5 The Impact of Pilot Size on Poisoning
The online learning of deep receivers is influenced by the training data, i.e., the pilot size $L_{\text{pilot}}$. Larger pilot sizes generally enhance performance, leading to more stable learning and reduced overfitting. To explore the impact of overfitting on the poisoning effect, Fig. 14 examines the influence of pilot size in a linear time-varying synthetic channel at an SNR of 14 dB. The results show that as $L_{\text{pilot}}$ increases, the effectiveness of poisoning attacks diminishes. Specifically, when comparing $L_{\text{pilot}}=200$ with $L_{\text{pilot}}=1000$, the poisoning effects on the ResNet detector, black-box DNN detector, DeepSIC, and Meta-DeepSIC are reduced by 0.34 dB, 0.74 dB, 0.14 dB, and 0.34 dB, respectively. However, Larger pilot sizes cannot alleviate the poisoning effects on meta-learning methods, and the poisoned Meta-DeepSIC still performs worse than DeepSIC.
@@ V-F Discussion of Results
The experimental results demonstrate that the poisoning attack method devised in this paper is effective in four distinct channel environments. Nevertheless, the precise impact of these attacks varies depending on the specific channel environment. In light of the aforementioned experimental results, the following inferences can be drawn.
@@
As illustrated in Fig.11, the poisoning effect observed in the nonlinear time-varying synthetic channel is markedly higher than in the other three cases. The performance of the deep receivers subjected to a poisoning attack is severely degraded and approaches failure in this channel environment. This suggests that the poisoning attack is capable of impeding the deep receiver’s ability to adapt to the rapid changes in the channel’s effects and to learn the nonlinear effects.
@@
Secondly, the experimental results for the linear time-varying synthetic channel and the COST 2100 channel corroborate the preceding conclusion from disparate vantages. In comparison to the synthetic channel, the tap coefficients of the COST 2100 channel demonstrate greater long-term variance, while exhibiting a relatively flat profile in the short term (for instance, between two blocks). As shown in Fig. 13(a), the impact of poisoning attacks consistently reduces receiver performance over an extended period in the COST 2100 channel, indicating a sustained but limited exacerbation. In contrast, in the linear time-varying synthetic channel undergoing significant changes over time, the poisoning attack has a more pronounced effect on the receiver’s performance. This indicates that the attack has a impact on the deep receiver’s capacity to track long-term channel alterations and a more pronounced disruptive effect on short-term rapid adaptation. Notably, as shown in Fig. 14, a larger pilot size can mitigate the adverse effects of poisoning attacks, but this means more spectrum resources are consumed.
@@
Finally, from the perspective of the deep receiver, Meta-DeepSIC, which incorporates a meta-learning approach, demonstrates optimal performance when utilizing limited pilot data, particularly in scenarios where channel variations occur rapidly. Furthermore, this learning capability demonstrates efficacy in nonlinear channel environments, while increased sensitivity to poisoned samples results in a more pronounced deterioration in performance than that observed for DeepSIC in fast-varying channels. In scenarios characterised by slow-varying or static channels, the performance of Meta-DeepSIC and DeepSIC, with or without poisoning attacks, is more consistent than in fast-varying channel conditions. This is corroborated by the data presented in Fig. 12 and Fig. 13.
@@
In conclusion, the attack method devised in this paper primarily impedes the receiver’s ability to learn rapid channel changes and non-linear effects in the short term, resulting in a decline in performance. The poisoning effect is particularly pronounced in the context of Meta-DeepSIC when combined with meta-learning techniques. This conclusion highlights the security risks associated with the design of wireless receivers using online and online meta-learning methods, particularly in environments characterised by rapid channel changes and non-linear effects, where the system is particularly susceptible to poisoning attacks.

### v3

@@ V-C Online Training Methods
The objective of this paper is to present an attack strategy, which attacks online adaptation of the deep receivers. Consequently, the focus is on the deep receiver’s online training, with the parameter configurations illustrated in Table I. The parameter settings for online training come from [15, 17]. Upon receipt of a data block, the deep receiver is only able to utilize a subset of the data block for training, like pilot symbols, specifically $L_{\text{pilot}}=200$. Then the deep reciver predicts the subsequent $L_{\text{info}}=50000$ symbols. In the experiment, a total of 100 data blocks were transmitted, with the transmitted data being QPSK modulated, i.e., the user-transmitted symbols $\mathbf{s}$ were mapped to the set $C=\left\{\left(\pm\frac{1}{\sqrt{2}},\pm\frac{1}{\sqrt{2}}\right)\right\}^{4}$. Furthermore, the deep receivers are trained using the adam optimizer. The training epochs are 300. The initial learning rate $\eta$ is set to 5 ×$\times 10^{-3}$ for ResNet detector, black-box DNN detector and DeepSIC. The meta learning rate $\eta_{\text{meta}}$ is set to 0.01 for Meta-DeepSIC. Additionally, online training is performed for different architectural receivers, which were implemented in the following two cases:
@@
    Online learning: Based on the adaptation from the previous data block, the deep receiver trains and updates the current model using the limited pilot symbols received in the current data block.
@@
    Online meta-learning: According to [20], [52], the meta-learning is employed to facilitate adaptation to the channel. Specifically, the pilot data from 5 data blocks is accumulated, and then meta-learning is performed on the this data to obtain the meta-learning weights for the deep receiver. Subsequently, the weights are then employed in an online learning process involving the pilot data of the current block.
Unless otherwise stated, the deep receivers are trained using online learning, including black-box DNN detector, ResNet detector, and DeepSIC. Only Meta-DeepSIC is trained using online meta-learning methods.
@@ V-D Attacker’s Configuration
The attack samples are designed based on the gradient of the surrogate model, with joint learning performed using the collected pilot dataset. The black-box DNN detector, presented in Section V-B, is employed as the surrogate model and contains three times the number of parameters compared to a single sub-network of DeepSIC. The joint learning parameters are based on [15]. In particular, the linear time-varying synthetic channel model is employed to generate the channel data for joint learning. Training data is generated under the condition of an SNR of 2 dB to 16 dB with an interval of 2 dB, denoted as $SNR_{\text{sur}}$. In this manner, $L_{\text{sur}}=5000$ training pilot symbols are generated at each SNR value, and the surrogate model is trained in accordance with this procedure.
Moreover, the poisoning samples are iteratively optimized using the PGD algorithm, following the settings in [44], where the adversarial poisoning samples are most effective under this specific parameter configuration. The iteration step size is set to $\gamma=0.01$, the iterations of PGD is $Q=250$, and the perturbation upper bound under the $p=2$ norm is $\epsilon=0.3$. The maximum and minimum values of the received symbol magnitude are denoted as $I_{max}$ and $I_{min}$, respectively, and $I_{max}=max\{\|\mathbf{y}\|\}=-I_{min}$. Fig. 8 shows the real and imaginary parts of the original and poisoned received symbols, while Fig. 9 presents these symbols in a constellation diagram.
@@ V-E Numerical Results under Four Channel Models
This section presents the experimental results evaluating the proposed poisoning attack across various channel models. The method’s effectiveness is tested on linear and nonlinear time-varying synthetic channels, linear static synthetic channels, and time-varying COST 2100 channels. For the black-box DNN detector, a white-box poisoning attack is launched against this architecture receiver since its architecture is the same as the surrogate model. In the case of the ResNet detector, transfer poisoning attacks are implemented in black-box scenarios. Since DeepSIC contains multiple sub-networks, direct gradient information cannot be used for designing poisoning samples. Therefore, a transfer-based poisoning attack is employed, anticipating that perturbations designed on the surrogate model will transfer to DeepSIC.
@@ V-E1 Linear Time-varying Synthesis Channel Results
Fig. 10(a) shows the results of 100 data transmission blocks, with the average cumulative SER calculated block-wise over five repetitions at an SNR of 14 dB. Meta-DeepSIC, combining model-based method with meta-learning, better captures the time-varying wireless channel characteristics in the absence of a poisoning attack, achieveing superior SER performance compared to the black-box DNN detector, ResNet detector and DeepSIC. However, Meta-DeepSIC is more susceptible to poisoning attacks than the DeepSIC. This aligns with the findings in [46]. Since this method targets the black-box DNN detector through a white-box attack approach, the black-box DNN detector performs the worst after poisoning. Meanwhile, the poisoning attack also significantly impacts the ResNet detector. Fig. 10(b) illustrates the results of the 100-block data transmission, with SER averaged over five repetitions under varying SNR conditions. The poisoning attack effectively degrades the performance of all four types of deep receivers at different SNRs. Specifically, the attack degrades SER by 0.68 dB for the ResNet detector, 0.5 dB for the black-box DNN detector, 0.67 dB for DeepSIC, and 0.91 dB for Meta-DeepSIC at SNR = 10 dB. As the SNR increases, the attack effect becomes more pronounced, with the SER deteriorations of the four receiver architectures reaching 0.68 dB, 0.43 dB, 1.40 dB, and 2.0 dB, respectively, at SNR = 15 dB.
@@ V-E2 Non-linear Time-varying Synthetic Channels Results
Fig. 11 evaluates the effectiveness of the proposed poisoning attack under a nonlinear time-varying synthetic channel, using the same experimental setup as in Fig. 10. The results indicate a stronger poisoning effect in the nonlinear channel compared to the linear one. Specifically, the attack degrades SER by 0.93 dB for the ResNet detector, 0.53 dB for the black-box DNN detector, 1.35 dB for DeepSIC, and 1.42 dB for Meta-DeepSIC at SNR = 10 dB. As the SNR increases, the attack effect becomes more pronounced, with the SER deteriorations of the four receiver architectures reaching 1.13 dB, 0.71 dB, 2.74 dB, and 3.09 dB, respectively, at SNR = 15 dB.
@@ V-E3 Linear Static Synthetic Channels Results
Fig. 12 evaluates the effectiveness of the proposed poisoning attack in a linear static synthetic channel, with the same setup as Fig. 10. The performance of DeepSIC and Meta-DeepSIC is nearly identical in this case. Furthermore, the lack of diverse data makes the data-driven detector unsuitable for adapting to the channel environment. For example, the black-box DNN detector struggles to adapt to the static channel, while the ResNet detector, which performed well in previous channels, performs the worst here, making the poisoning effect negligible. In particular, at SNR = 10 dB, the poisoning attack degrades the SER by 0.08 dB for the black-box detector, 0.71 dB for DeepSIC, and 0.72 dB for Meta-DeepSIC. As the SNR increases, the impact intensifies, with SER degradations of 0.06 dB, 1.05 dB, and 1.07 dB at SNR = 15 dB, respectively.
@@ V-E4 Time-varying COST 2100 Channels Results
The effectiveness of the proposed poisoning attack is evaluated under the time-varying COST 2100 channel, as shown in Fig. 13. The experimental configuration is identical to that of the linear time-varying synthetic channel. Simultaneously, the surrogate model is trained on the joint channel data based on the time-varying linear channel model. As Fig. 13 illustrates, the poisoning attack proves effective across receivers with varying architectural and online training scenarios, while also adapting to different channel environments. At SNR = 10 dB, the attack degrades SER performance by 0.12 dB for the ResNet detector, 0.25 dB for the black-box DNN detector, 0.61 dB for DeepSIC, and 0.78 dB for Meta-DeepSIC. At SNR = 15 dB, the attack degrades SER performance by 0.04 dB for the ResNet detector, 0.23 dB for the black-box DNN detector, 2.11 dB for DeepSIC, and 1.36 dB for Meta-DeepSIC.
@@ V-E5 The Impact of Pilot Size on Poisoning
The online learning of deep receivers is influenced by the training data, i.e., the pilot size $L_{\text{pilot}}$. Larger pilot sizes generally enhance performance, leading to more stable learning and reduced overfitting. To explore the impact of overfitting on the poisoning effect, Fig. 14 examines the influence of pilot size in a linear time-varying synthetic channel at SNR = 14 dB. The results show that as $L_{\text{pilot}}$ increases, the effectiveness of poisoning attacks diminishes. Specifically, when comparing $L_{\text{pilot}}=200$ with $L_{\text{pilot}}=1000$, the poisoning effects on the ResNet detector, black-box DNN detector, DeepSIC, and Meta-DeepSIC are reduced by 0.34 dB, 0.74 dB, 0.14 dB, and 0.34 dB, respectively. However, Larger pilot sizes cannot alleviate the poisoning effects on meta-learning methods, and the poisoned Meta-DeepSIC still performs worse than poisoned DeepSIC.
@@ V-F Discussion of Results
The experimental results demonstrate that the proposed poisoning attack is effective across four different channel environments. Nevertheless, the precise impact of these attacks varies depending on the specific channel environment. Firstly, as illustrated in Fig.11, the poisoning effect observed in the nonlinear time-varying synthetic channel is markedly higher than in the other environments. The performance of the deep receivers subjected to a poisoning attack is severely degraded and approaches failure in this channel environment. This suggests that the poisoning attack is capable of impeding the deep receiver’s ability to adapt to the rapid changes in the channel’s effects and to learn the nonlinear effects.
@@
Secondly, the results for the linear time-varying synthetic channel and the COST 2100 channel support this conclusion from different perspectives. In comparison to the synthetic channel, the tap coefficients of the COST 2100 channel demonstrate greater long-term variance, while exhibiting a relatively flat profile in the short term (e.g., between two blocks). As illustrated in Fig. 13(a), the impact of poisoning attacks gradually reduces receiver performance in the COST 2100 channel, indicating a sustained but limited exacerbation. In contrast, in the linear time-varying synthetic channel undergoing significant changes over time, the poisoning attack has a more pronounced effect on the receiver’s performance. This indicates that the attack has a impact on the deep receiver’s capacity to track long-term channel alterations and a more pronounced disruptive effect on short-term rapid adaptation. Notably, as shown in Fig. 14, a larger pilot size can mitigate the adverse effects of poisoning attacks, but this means more spectrum resources are consumed.
@@
Finally, Meta-DeepSIC, which incorporates a meta-learning approach, demonstrates optimal performance with limited pilot data, particularly in fast-varying channels. Additionally, this learning capability is effective in nonlinear environments, but its increased sensitivity to poisoned samples leads to more significant performance deterioration compared to DeepSIC in fast-varying channels. In slow-varying or static channels, the performance of both Meta-DeepSIC and DeepSIC, with or without poisoning attacks, is more stable than in fast-varying channels, as shown in Fig. 12 and Fig. 13.
@@
In conclusion, the proposed attack primarily impedes the receiver’s ability to learn rapid channel changes and non-linear effects in the short term, leading to performance degradation. Meanwhile, when deep receivers are combined with meta-learning, the poisoning effect is particularly pronounced (e.g., Meta-DeepSIC). This conclusion highlights the security risks associated with designing wireless receivers using online and online meta-learning methods, particularly in environments characterised by rapid channel changes and non-linear effects, where the system is particularly susceptible to poisoning attacks.